There are two competing brands of tooth paste in a market the brands are colget and akuafresh. A survey was undertaken to determine the extent consumer switch over from one brand to another every quarter. The survey has identified that every quarter 1/5 of 45% of akuafresh and 3/4 of 60% customers of colget in the previous quarter become customer of akuafresh. Moreover it was identified that the two brands have constant customer of 10,000 and no one stops purchasing one or the other brand currently the number of customers of colget are 6000 persons
A,what will be the number of customers of colget and akuafresh after two transitions?
B, find the steady stake market share
1
Expert's answer
2021-09-14T06:05:29-0400

"Colgate: 6000 \\
Aquafresh: 4000\\
First: \\
Total: 10000\\
 4000×45\%=1800 \\
Now, \frac{1}{3} \ rd \ of \ 1800=600\\
Also, 6000×60\%=3600\\
\therefore C:6000-3600+600=3000 \\
A:4000+3600-600=7000 \\

Second: \\
7000×45\%×\frac{1}{3}=1050 \\
Also, 3000×60\%=1800\\
C:3000-1800+1050=2250 \\
A:10000-2250=7750"